OpenStudy (anonymous):

I need someone to check this question for me. I solved the system using elimination 3x-9=3 6x-3y=-24 3x-9y=3 18x-9y = -72 = -15x-0=75 x= -5 6x-3y=-24 6(-5)-3y=-24 -30-3y=-24 y=-2

OpenStudy (ajprincess):

Perfect.:)

OpenStudy (unklerhaukus):

\[3x-9y=3\]\[6x-3y=-24\] is \((x,y)=(-5,-2)\) a solution? does \(3(-5)-9(-2)=3\)? and \[6(-5)-3(-2)=-24\]?

OpenStudy (anonymous):

Thank you for checking it for me :)

OpenStudy (ajprincess):

u r welcome.